OpenStudy (31356):

Factoring Using the Distributive Property 5a^2-15 Show your steps to find the GCF MEDAL REWARDED!!

OpenStudy (31356):

@satellite73

OpenStudy (31356):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do 5 and 15 have in common?

jimthompson5910 (jim_thompson5910):

in terms of factors

OpenStudy (31356):

5

jimthompson5910 (jim_thompson5910):

so factor out the GCF to get ???

jimthompson5910 (jim_thompson5910):

5a^2 - 15 = 5( .... ) what goes in the parenthesis?

OpenStudy (31356):

5? I don't know

OpenStudy (31356):

I just learned this today

jimthompson5910 (jim_thompson5910):

factor 15 into 2 numbers where 5 is one factor

OpenStudy (31356):

In class

jimthompson5910 (jim_thompson5910):

15 = 5*____

OpenStudy (31356):

3

jimthompson5910 (jim_thompson5910):

15 = 5*3, good

jimthompson5910 (jim_thompson5910):

so 5a^2 - 15 5a^2 - 5*3 5(a^2 - 3)

jimthompson5910 (jim_thompson5910):

for that last step, I'm using the distributive property a*(b+c) = a*b + a*c

OpenStudy (31356):

Thanks!

OpenStudy (31356):

So this is the answer right? 5(a^2 - 3)

jimthompson5910 (jim_thompson5910):

Notice as a check, we can distribute that 5 back through to get 5(a^2 - 3) 5(a^2) + 5(-3) 5a^2 - 15 which is the original expression. So the answer is confirmed

jimthompson5910 (jim_thompson5910):

Yes, it's 5(a^2 - 3)

OpenStudy (31356):

Ok, thanks so much!

jimthompson5910 (jim_thompson5910):

sure thing